### v3

# Solutions for $k$-generalized Fibonacci numbers using Fuss-Catalan numbers

Abstract
We present new expressions for the $k$-generalized Fibonacci numbers, say $F_{k}(n)$. They satisfy the recurrence $F_{k}(n)=F_{k}(n-1)+\dots+F_{k}(n-k)$. Explicit expressions for the roots of the auxiliary (or characteristic) polynomial are presented, using Fuss-Catalan numbers. Properties of the roots are enumerated. We quantify the accuracy of asymptotic approximations for $F_{k}(n)$ for $n\gg 1$. Our results subsume and extend some results published by previous authors. We also comment on the use of multinomial sums for the $k$-generalized Fibonacci numbers and related sequences. We also employ the generating function to express $F_{k}(n)$ as a concise (non-nested) sum of binomial coefficients for arbitrary $k$. Finally, we present a basis (or ‘fundamental solutions’) to solve the above recurrence for arbitrary initial conditions.
keywords:
Generalized Fibonacci numbers , Fuss-Catalan numbers , recurrences , Vandermonde systems
MSC:
[2020] 11B39 , 11B37 , 39A06
journal: (internal report CC24-8)
1 Introduction
We treat the so-called ‘$k$-generalized’ Fibonacci numbers, say $F_{k}(n)$. They satisfy the recurrence
$F_{k}(n)=F_{k}(n-1)+\dots+F_{k}(n-k)\,.$ (1.1)
The usual Fibonacci numbers are the special case $k=2$. Eq. (1.1) is a homogeneous linear difference equation and its characteristic equation is
$x^{k}-\sum_{j=0}^{k-1}x^{j}=0\,.$ (1.2)
The corresponding auxiliary (or characteristic) polynomial is $A(x)=x^{k}-x^{k-1}-\dots-1$. Several authors [1, 2, 3] have shown that the roots of $A(x)$ are simple. Let us denote the roots of $A(x)$ by $\zeta_{\ell}$, $\ell=0,\dots,k-1$. Then we express $F_{k}(n)$ as a weighted sum of powers of the roots as follows (i.e. a Binet-style formula)
$F_{k}(n)=\sum_{\ell=0}^{k-1}c_{\ell}\zeta_{\ell}^{n}\,.$ (1.3)
Here $\{c_{0},\dots,c_{k-1}\}$ is a set of coefficients which do not depend on $n$, but they do depend on $k$ and the initial conditions, i.e. the values of $F_{k}(0)$ through $F_{k}(k-1)$. It is standard to employ the initial conditions $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. We shall begin with this, but later we shall treat alternative initial conditions, by Spickerman and Joyner [4]. In Sec. 8, we present a complete set of ‘fundamental solutions’ for the recurrence, i.e. a basis which can be used to solve for arbitrary initial conditions.
Our main result is to present explicit expressions for the roots $\zeta_{j}$, in terms of so-called ‘Fuss-Catalan numbers’ (see the text by Graham, Knuth and Patashnik [5] for relevant definitions, formulas and identities). We also employ the generating function to express $F_{k}(n)$ as a concise (non-nested) sum of binomial coefficients for arbitrary $k$ (see Sec. 7). Finally, in Sec. 8 we present a basis (or ‘fundamental solutions’) to solve the above recurrence for arbitrary initial conditions. The connection to Fuss-Catalan numbers is somewhat surprising. Dilworth and Mane [6] (hereafter ‘DM’) published a paper on success run waiting times of Bernoulli trials. The recurrence for the probability mass function $f_{k}(n)$ is ([6] eq. (1)), with $p\in(0,1)$, $q=1-p$ and $k\geq 2$,
$f_{k}(n)=qf_{k}(n-1)+pqf_{k}(n-2)+p^{2}qf_{k}(n-3)+\cdots+p^{k-1}qf_{k}(n-k)\,.$ (1.4)
Set $p=q=\frac{1}{2}$ and $f_{k}(n)=F_{k}(n)/2^{n}$; then elementary manipulations yield eq. (1.1). Hence the solution for the $k$-generalized Fibonacci numbers $F_{k}(n)$ equals $2^{n}$ times the probability mass function $f_{k}(n)$ of success run waiting times of Bernoulli trials, with $p=q=\frac{1}{2}$. The auxiliary polynomial associated with eq. (1.4) is $\mathcal{A}_{p}(x)=x^{k}-qx^{k-1}-pqx^{k-2}-\cdots-p^{k-1}q$. Barry and Lo Bello [7] showed that the roots of $\mathcal{A}_{p}(x)$ are simple, which goes to show that there are fruitful connections between different fields of research. Let us denote the roots of $\mathcal{A}_{p}(x)$ by $\lambda_{\ell}$, $\ell=0,\dots,k-1$. Then $\zeta_{\ell}=2\lambda_{\ell}$, with $p=q=\frac{1}{2}$.
It is known that $\mathcal{A}_{p}(x)$ has a unique positive root, which has a larger magnitude than all the other roots [8]. We denote it by $\lambda_{0}$ below. It is also known that if $k$ is even then there is a unique negative root. For all $k$, the remaining roots come in complex conjugate pairs [6]. Wolfram [3] also proved these results for the case of $k$-generalized Fibonacci numbers. DM employed the terms “primary root” for $\lambda_{0}$ and “secondary roots” for the other $\lambda_{j}$, $j=1,\dots,k-1$.
DM obtained expressions for the roots $\lambda_{j}$ as power series, whose coefficients are Fuss-Catalan numbers ([6] Theorem 1). They identified three cases, viz. (i) $p<k/(k+1)$, (ii) $p=k/(k+1)$ and (iii) $p>k/(k+1)$. We require only the first case, because we fix $p=\frac{1}{2}$ and $k/(k+1)>\frac{1}{2}$ for all $k>1$. First define, for $n\geq 1$,
$b_{n}=\frac{\Gamma(n-1+n/k)}{n!\Gamma(n/k)}\,.$ (1.5)
The Fuss-Catalan numbers are defined as follows ([6] eq. (9))
$A_{m}(\nu,r)=\frac{r}{\Gamma(m+1)}\frac{\Gamma(m\nu+r)}{\Gamma(m(\nu-1)+r+1)}\,.$ (1.6)
Then $b_{n}$ is a Fuss-Catalan number ([6] eq. (10))
$b_{n}=-A_{n}(1+1/k,-1)\,.$ (1.7)
The primary root $\zeta_{0}$ is given by ([6] eq. (8))
$\zeta_{0}=2-2k\sum_{m=1}^{\infty}b_{mk}\,(p^{k}q)^{m}=2-2k\sum_{m=1}^{\infty}\frac{b_{mk}}{2^{m(k+1)}}\,.$ (1.8)
The secondary roots $\zeta_{j}$, $j=1,\dots,k-1$ are given by ([6] eq. (7))
$\zeta_{j}=2\sum_{n=1}^{\infty}b_{n}\,(e^{2\pi ij/k}pq^{1/k})^{n}=2\sum_{n=1}^{\infty}b_{n}\,\biggl{(}\frac{e^{2\pi ij/k}}{2^{(k+1)/k}}\biggr{)}^{n}\,.$ (1.9)
Wolfram ([3] Theorem 3.9) derived the following sum for the principal root. The root was expressed as $\zeta_{0}=2(1-\varepsilon_{k})$ and
$\varepsilon_{k}=\sum_{i\geq 1}\binom{(k+1)i-2}{i-1}\frac{1}{i2^{(k+1)i}}\,.$ (1.10)
This agrees with eq. (1.8); the proof is as follows.
Proof.
From eq. (1.5), note that
$b_{mk}=\frac{\Gamma(mk-1+m)}{(mk)!\Gamma(m)}=\frac{(mk+m-2)!}{(mk!)(m-1)!}=\frac{1}{mk}\binom{(k+1)m-2}{m-1}\,.$ (1.11)
Next using eq. (1.8), we deduce
$\begin{split}\varepsilon_{k}&=k\sum_{m=1}^{\infty}\frac{b_{mk}}{2^{(k+1)m}}\\ &=k\sum_{m=1}^{\infty}\frac{1}{mk}\binom{(k+1)m-2}{m-1}\frac{1}{2^{(k+1)m}}\\ &=\sum_{m=1}^{\infty}\binom{(k+1)m-2}{m-1}\frac{1}{m2^{(k+1)m}}\,.\end{split}$ (1.12)
This equals eq. (1.10), substitute $i$ for $m$. ∎
However, Wolfram did not publish expressions for the secondary roots. Spickerman and Joyner ([4] Sec. 3) computed the roots numerically (to $4$ decimal places) for $2\leq k\leq 10$. Using eqs. (1.8) and (1.9), we confirm their numbers are correct.
It remains now to derive expressions for $F_{k}(n)$ for specific initial conditions, and also to derive asymptotic expressions for $n\gg 1$ (in which case the solution is dominated by the contribution from the principal root). To do so, it will be helpful below to work with the polynomial $B(x)=(x-1)A(x)$ for the following reason. Note that
$\begin{split}A(x)&=x^{k}-x^{k-1}-\dots-1\\ &=x^{k}-\frac{x^{k}-1}{x-1}\\ &=\frac{x^{k+1}-2x^{k}+1}{x-1}\,.\end{split}$ (1.13)
Hence $B(x)=x^{k+1}-2x^{k}+1$ is a trinomial, which is easier to manipulate. However, it contains an extraneous root $x=1$. Note that $A(1)=1-k\neq 0$, i.e. $x=1$ is not a root of $A(x)$, hence $B(x)$ has no repeated roots.
Remark 1.1.
Gabai ([9] Theorem 1) published the following recurrence, which can also be deduced from $B(x)$ above
$F_{k}(n+1)=2F_{k}(n)-F_{k}(n-k)\,.$ (1.14)
For all $k\geq 2$, there are only two terms on the right-hand side. Initial values must be supplied for $n=0,\dots,k$. Eq. (1.14) has the extraneous solution $F_{k}(n)=1$ for all $n$; the choice of initial values will exclude this case.
2 Properties of roots
We may read off several properties of the roots from DM [6]. First, the principal root lies in the interval ([6] Props. 8 and 11)
$2-\frac{2}{k+1}<\zeta_{0}<2-\frac{1}{2^{k}}\,.$ (2.1)
Wolfram ([3] Lemma 3.6) derived the bounds
$2(1-2^{-k})<\zeta_{0}<2\,.$ (2.2)
Eq. (2.1) has a tighter upper bound and eq. (2.2) has a tighter lower bound. Combining them, we obtain the improved bounds
$2-\frac{1}{2^{k-1}}<\zeta_{0}<2-\frac{1}{2^{k}}\,.$ (2.3)
Next, DM ([6] Prop. 14) showed that the secondary roots satisfy $|\lambda_{j}|<p$. Setting $p=\frac{1}{2}$, this implies $|\zeta_{j}|<1$ for $j=1,\dots,k-1$. Wolfram ([3] Lemma 3.6) obtained a tighter lower bound
$3^{-1/k}<|\zeta_{j}|<1\,.$ (2.4)
This corrects a misprint in [3], where the lower bound is given as $3^{-k}$ (but the value $3^{-1/k}$ is stated correctly in the proof of Lemma 3.6 in [3]). Next, the relative magnitudes of distinct roots are as follows ([6] Prop. 16)
  1. 1.
    $|\zeta_{j_{1}}|=|\zeta_{j_{2}}|$ if and only if $\zeta_{j_{1}}=\bar{\zeta}_{j_{2}}$.
  2. 2.
    $\Re(\zeta_{j_{1}})<\Re(\zeta_{j_{2}})$ if and only if $|\zeta_{j_{1}}|<|\zeta_{j_{2}}|$.
  3. 3.
    For even $k$, the negative real root has the most negative real part, hence the smallest amplitude of all the roots.
Spickerman and Joyner ([4] Sec. 3) computed the roots numerically for $2\leq k\leq 10$ and their numbers satisfy the above relations.
Distribution of the roots in the complex plane. Observe the following limits as $k\to\infty$: (i) for the principal root, $\zeta_{0}\to 2^{-}$ (from eq. (2.3)) and (ii) for all the secondary roots, $|\zeta_{j}|\to 1^{-}$ (from eq. (2.4)). Hence all the secondary roots approach the unit circle (from the inside), but the principal root approaches $2$ (from below). A graph of the roots in the complex plane is plotted in Fig. 1, for $k=40$. For $k\gg 1$, the secondary roots are also distributed approximately uniformly around the unit circle. A graph of the arguments of the secondary roots, i.e. $\arg(\zeta_{j})/(2\pi)$, is plotted in Fig. 2, for $k=40$. The points lie close to a straight line.
3 Vandermonde
First recall eq. (1.3). We set up a Vandermonde system of equations to solve for the coefficients $c_{\ell}$ as follows
$\begin{split}\begin{pmatrix}1&1&\dots&1\\ \zeta_{0}&\zeta_{1}&\dots&\zeta_{k-1}\\ \zeta_{0}^{2}&\zeta_{1}^{2}&\dots&\zeta_{k-1}^{2}\\ \vdots&&&\vdots\\ \zeta_{0}^{k-1}&\zeta_{1}^{k-1}&\dots&\zeta_{k-1}^{k-1}\end{pmatrix}\begin{pmatrix}c_{0}\\ c_{1}\\ c_{2}\\ \vdots\\ c_{k-1}\end{pmatrix}=\begin{pmatrix}F_{k}(0)\\ F_{k}(1)\\ F_{k}(2)\\ \vdots\\ F_{k}(k-1)\end{pmatrix}\,.\end{split}$ (3.1)
We begin with the initial conditions $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. (We treat the initial condition by Spickerman and Joyner [4] in Sec. 5.) The solution is
$c_{\ell}=\frac{1}{\prod_{j\neq\ell}(\zeta_{\ell}-\zeta_{j})}\,.$ (3.2)
Note that
$B(x)=(x-1)\prod_{j=0}^{k-1}(x-\zeta_{j})\,.$ (3.3)
Hence
$B^{\prime}(x)=\frac{B(x)}{x-1}+\sum_{j=0}^{k-1}\frac{B(x)}{x-\zeta_{j}}\,.$ (3.4)
Then, noting that $B(\zeta_{\ell})=0$,
$B^{\prime}(\zeta_{\ell})=(\zeta_{\ell}-1)\prod_{j\neq\ell}(\zeta_{\ell}-\zeta_{j})\,.$ (3.5)
Also $B^{\prime}(x)=(k+1)x^{k}-2kx^{k-1}$, thus $B^{\prime}(\zeta_{\ell})=((k+1)\zeta_{\ell}-2k)\zeta_{\ell}^{k-1}$. Hence
$\begin{split}c_{\ell}&=\frac{1}{\prod_{j\neq\ell}(\zeta_{\ell}-\zeta_{j})}=\frac{\zeta_{\ell}-1}{B^{\prime}(\zeta^{\ell})}=\frac{\zeta_{\ell}-1}{((k+1)\zeta_{\ell}-2k)\zeta_{\ell}^{k-1}}\,.\end{split}$ (3.6)
Hence the solution for $F_{k}(n)$ is
$F_{k}(n)=\sum_{\ell=0}^{k-1}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\zeta_{\ell}^{n-k+1}\,.$ (3.7)
This is our expression for the $k$-generalized Fibonacci numbers, as a sum over powers of the roots of the auxiliary polynomial. The expressions for special cases such as the tribonacci numbers ($k=3$) and tetranacci numbers ($k=4$) are known in radicals. However eq. (3.7) presents a unified formula for all $k\geq 2$. The use of Fuss-Catalan numbers furnishes a systematic formula to compute the roots to arbitrary accuracy.
Remark 3.1.
An expression equivalent to eq. (3.7) was published by Dresden and Du ([10] Theorem 1), where the roots were denoted by $\alpha_{j}$, $j=1,\dots,k$. However, they did not exhibit expressions for the roots. They denoted the $k$-generalized Fibonacci numbers by $F_{n}^{(k)}$, with the recurrence $F_{n}^{(k)}=F_{n-1}^{(k)}+\dots+F_{n-k}^{(k)}$, but with the initial conditions $F_{n}^{(k)}=0$ for $n<1$ and $F_{n}^{(k)}=1$ for $n=1$. The relation to our notation is $F_{k}(n)=F_{n-k+2}^{(k)}$.
4 Asymptotic solution
For fixed $k$ and $n\gg k$, the asymptotic solution is given by retaining only the principal root $\zeta_{0}$ in eq. (3.7). The details of the analysis are given in DM ([6] Corollary 4). Use $B(\zeta_{0})=0$, i.e. $\zeta_{0}^{k}(\zeta_{0}-2)+1=0$ to deduce $\zeta_{0}^{-k}=2-\zeta_{0}$. Then (c.f. [6], eq. (16), case $p\neq k/(k+1)$)
$\begin{split}F_{k}(n)&\asymp\frac{\zeta_{0}-1}{(k+1)\zeta_{0}-2k}\zeta_{0}^{n-k+1}\\ &=\frac{(2-\zeta_{0})(\zeta_{0}-1)}{(k+1)\zeta_{0}-2k}\zeta_{0}^{n+1}\,.\end{split}$ (4.1)
Following DM, we demand that for fixed $\varepsilon>0$, the magnitude of the contribution to $F_{k}(n)$ in eq. (3.7) from all the secondary roots combined is less than $\varepsilon$ times the contribution from the principal root. This is achieved if $n\geq N(p,k)$, where ([6], eq. (17))
$N=k+\frac{\ln(\varepsilon/(k-1))}{\ln\kappa}\,.$ (4.2)
In our case $\kappa$ is given by ([6], eq. (18), case $1/(k+1)<p<k/(k+1)$)
[…]
Here ([6], eq. (19))
$\delta=\frac{1}{(k+1)^{1/k}}\biggl{[}1+\frac{1}{2^{(k+1)/k}k}-\sqrt{1+\frac{1}{2^{2(k+1)/k}k^{2}}+\frac{\cos(2\pi/k)}{2^{1/k}k}}\biggr{]}\,.$ (4.4)
Then $\delta\to 0$ as $k\to\infty$, hence the expression for $\kappa$ simplifies to $\kappa=(k+1)/(2k)$ and we obtain the bound
$N=k+\frac{\ln(\varepsilon/(k-1))}{\ln((k+1)/(2k))}\,.$ (4.5)
Remark 4.1.
The above analysis was leveraged from [6]. For the specific application to $k$-generalized Fibonacci numbers, Dresden and Du [10] derived a simpler expression. They showed ([10] Theorem 2) that the contribution of all the secondary roots is less than $\frac{1}{2}$ in magnitude, and $F_{k}(n)$ can be obtained from the principal root $\zeta_{0}$ alone as follows. We employ our notation in eq. (3.7), whence
$F_{k}(n)=\left\lfloor\biggl{(}\frac{\zeta_{0}-1}{(k+1)\zeta_{0}-2k}\zeta_{0}^{n-k+1}+\frac{1}{2}\biggr{)}\right\rfloor\,.$ (4.6)
5 Alternative initial conditions: Spickerman and Joyner
Spickerman and Joyner [4] solved the recurrence eq. (1.1) as follows. First to avoid confusions of notation, we follow them and write $u_{n}$ in place of $F_{k}(n)$. They employed the initial conditions $(u_{0},\dots,u_{k-1})=(1,1,2,4,\dots,2^{k-2})$, i.e. $u_{0}=1$ and $u_{n}=2^{n-1}$ for $11\leq n\leq k-$ 1. Hence the next term is $u_{k}=1+1+2+4+\dots+2^{k-2}=2^{k-1}$. Let us change this slightly and shift the indices down by one (and discard $u_{0}$) and say instead $(v_{0},\dots,v_{k-1})=(1,2,4,\dots,2^{k-1})$, i.e. $v_{n}=2^{n}$ for $10\leq n\leq k-$ 1. Let us generalize this even more and replace $2$ by $\mu$, and write $w_{n}=\mu^{n}$ for $10\leq n\leq k-$ 1. Then the solution for $c_{\ell}$ in the Vandermonde matrix equations in Sec. 3 is easy; it is the Lagrange polynomial $c_{\ell}=L_{\ell}(\mu)$. The Lagrange polynomial $L_{\ell}(x)$ is
$L_{\ell}(x)=\frac{(x-\zeta_{0})\dots(x-\zeta_{\ell-1})(x-\zeta_{\ell+1})\dots(x-\zeta_{k-1})}{(\zeta_{\ell}-\zeta_{0})\dots(\zeta_{\ell}-\zeta_{\ell-1})(\zeta_{\ell}-\zeta_{\ell+1})\dots(\zeta_{\ell}-\zeta_{k-1})}\,.$ (5.1)
This can be obtained from the auxiliary polynomial, written in the form (see eq. (1.13)) $A(x)=(x^{k+1}-2x^{k}+1)/(x-1)$. The denominator of $L_{\ell}(x)$ is $D_{\ell}=A^{\prime}(\zeta_{\ell})$ and does not depend on $x$. The numerator of $L_{\ell}(\mu)$ is $N_{\ell}(\mu)=A(\mu)/(\mu-\zeta_{\ell})$, provided $\mu$ is not a root of $A(x)$. For now, we assume $\mu$ is not a root of $A(x)$. Then the numerator is
$N_{\ell}(\mu)=\frac{\mu^{k+1}-2\mu^{k}+1}{(\mu-1)(\mu-\zeta_{\ell})}\,.$ (5.2)
The derivative of $A(x)$ is
$A^{\prime}(x)=\frac{(k+1)x^{k}-2kx^{k-1}}{x-1}-\frac{x^{k+1}-2x^{k}+1}{(x-1)^{2}}\,.$ (5.3)
The denominator is given by (note that the second term in $A^{\prime}(x)$ vanishes for $x=\zeta_{\ell}$)
$D_{\ell}=\frac{(k+1)\zeta_{\ell}^{k}-2k\zeta_{\ell}^{k-1}}{\zeta_{\ell}-1}=\frac{(k+1)\zeta_{\ell}-2k}{\zeta_{\ell}-1}\zeta_{\ell}^{k-1}\,.$ (5.4)
Hence the coefficient $c_{\ell}$ is
$c_{\ell}=\frac{N_{\ell}(\mu)}{D_{\ell}}=\frac{\mu^{k+1}-2\mu^{k}+1}{(\mu-1)(\mu-\zeta_{\ell})}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}\,.$ (5.5)
For $\mu=1$, set $\mu=1+\varepsilon$ and take the limit $\varepsilon\to 0$ to obtain
$\lim_{\mu\to 1}c_{\ell}=\frac{1-k}{(1-\zeta_{\ell})}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}=\frac{k-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}\,.$ (5.6)
What if $\mu$ is a root of $A(x)$? Actually this is easy. If $\mu=\zeta_{i}$ then by construction of the Lagrange polynomials, $L_{i}(\zeta_{i})=1$ and $L_{\ell}(\zeta_{i})=0$ for $i\neq\ell$. Hence $c_{i}=1$ and $c_{\ell}=0$ for $i\neq\ell$, i.e. $c_{\ell}=\delta_{i\ell}$.
Spickerman and Joyner [4] treated only the case $\mu=2$. We know that $A(2)=1$, hence $2$ is not a root of $A(x)$, whence
$\begin{split}c_{\ell}&=\frac{2^{k+1}-2^{k+1}+1}{(2-1)(2-\zeta_{\ell})}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}\\ &=\frac{1}{1/\zeta_{\ell}^{k}}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\frac{1}{\zeta_{\ell}^{k-1}}\\ &=\zeta_{\ell}\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\end{split}$ (5.7)
Let us compare this to Spickerman and Joyner’s solution [4]. Because of the index shift by $1$, the relation of $w$ to $u$ is $w_{n}(\mu=2)=u_{n+1}$. Write
$u_{n}=\sum_{j=0}^{k-1}d_{\ell}\zeta_{k}^{n}\,.$ (5.8)
Our claim is $c_{\ell}\zeta_{\ell}^{n}=d_{\ell}\zeta_{\ell}^{n+1}$, or $c_{\ell}=d_{\ell}\zeta_{\ell}$. Note that Spickerman and Joyner wrote $\alpha_{j}$, whereas our notation is $\zeta_{\ell}$. From Spickerman and Joyner (unnumbered, second last line on p. 328 of [4])
$d_{\ell}=\zeta_{\ell}^{k}\frac{\zeta_{\ell}-1}{2\zeta_{\ell}^{k}-(k+1)}\,.$ (5.9)
Equate and compare
$\begin{split}\frac{c_{\ell}-d_{\ell}\zeta_{\ell}}{\zeta_{\ell}(\zeta_{\ell}-1)}&=\frac{1}{(k+1)\zeta_{\ell}-2k}-\frac{\zeta_{\ell}^{k}}{2\zeta_{\ell}^{k}-(k+1)}\\ &\propto(2\zeta_{\ell}^{k}-(k+1))-\zeta_{\ell}^{k}((k+1)\zeta_{\ell}-2k)\\ &=2\zeta_{\ell}^{k}-(k+1)-(k+1)\zeta_{\ell}^{k+1}+2k\zeta_{\ell}^{k}\\ &=2\zeta_{\ell}^{k}-2(k+1)\zeta_{\ell}^{k}+2k\zeta_{\ell}^{k}\\ &=0\,.\end{split}$ (5.10)
Our claim is validated.
Note that although Spickerman and Joyner [4] gave a correct formal expression for the coefficient $d_{\ell}$, they did not offer an explicit expression for the roots of the auxiliary polynomial. Instead they solved the auxiliary equation numerically for $k=2,\dots,10$ and tabulated the (approximate) numerical values of the roots ([4] Sec. 3). As stated above, our expressions using Fuss-Catalan numbers match their numbers.
6 Multinomial sum
6.1 General remarks
We digress to consider alternative expressions using multinomial sums. Unlike the previous sections, this is an independent statement of known results. It is included for completeness of the exposition.
6.2 $k$-generalized Fibonacci
Lee, Lee, Kim and Shin [11] derived the Vandermonde set of equations, and published the formal solution, but did not find the roots explicitly. Their last formula (unnumbered) expresses the $k$-generalized Fibonacci numbers using multinomial sums. In terms of our notation, $F_{k}(0)=\dots=F_{k}(k-2)=0$, $F_{k}(k-1)=1$ and for $n\geq 0$
$\begin{split}F_{k}(n+k-1)&=\sum_{i_{1}+2i_{2}+\dots+ki_{k}=n}\frac{(i_{1}+i_{2}+\dots+i_{k})!}{i_{1}!i_{2}!\dots i_{k}!}\\ &=\sum_{i_{1}+2i_{2}+\dots+ki_{k}=n}\binom{i_{1}+i_{2}+\dots+i_{k}}{i_{1},i_{2},\dots,i_{k}}\,.\end{split}$ (6.1)
The proof is given by employing the generating function for the $k$-generalized Fibonacci numbers
$G(z)=\frac{z^{k-1}}{1-z-z^{2}-\dots-z^{k}}\,.$ (6.2)
Expand the denominator using the multinomial theorem and collect terms to obtain the coefficient of $z^{n+k-1}$.
6.3 More general recurrence
We can generalize the recurrence and multinomial sum as follows. Suppose the recurrence has coefficients $\{\beta_{1},\dots,\beta_{k}\}$, so the recurrence is, omitting too many labels which clutter the notation,
$F(n)=\beta_{1}F(n-1)+\cdots+\beta_{k}F(n-k)\,.$ (6.3)
The generating function is
$G(z;\beta_{1},\dots,\beta_{k})=\frac{z^{k-1}}{1-\beta_{1}z-\dots-\beta_{k}z^{k}}\,.$ (6.4)
Then the multinomial sum is, for $n\geq 0$,
$F(n+k-1)=\sum_{i_{1}+2i_{2}+\dots+ki_{k}=n}\binom{i_{1}+i_{2}+\dots+i_{k}}{i_{1},i_{2},\dots,i_{k}}\,\beta_{1}^{i_{1}}\dots\beta_{k}^{i_{k}}\,.$ (6.5)
Suitable initial values must be specified for $F(0)$ through $F(k-2)$. Note also that the sum in eq. (6.5) is normalized to yield $F(k-1)=1$.
6.4 Narayana sequence
The Narayana sequence is given by the initial values $N(0)=N(1)=N(2)=1$ and for $n\geq 3$ by the recurrence
$N(n)=N(n-1)+N(n-3)\,.$ (6.6)
Hence we set $k=3$ and $(\beta_{1},\beta_{2},\beta_{3})=(1,0,1)$ in Sec. 6.3. Because $\beta_{2}=0$, we exclude all terms in $i_{2}$ in eq. (6.5) with $i_{2}>0$. This yields, using $i_{1}=n-3i_{3}$,
$F(n+2)=\sum_{i_{1}+3i_{3}=n}\binom{i_{1}+i_{3}}{i_{1},i_{3}}=\sum_{i_{3}=0}^{\lfloor(n/3)\rfloor}\binom{n-2i_{3}}{i_{3}}\,.$ (6.7)
Set the initial values $F(0)=F(1)=1$, then $F(n)=N(n)$. The first few terms are $1,1,1,2,3,4,6,9,13,19,\dots$.
6.5 Padovan sequence
The Padovan sequence is given by the initial values $P(0)=P(1)=P(2)=1$ and for $n\geq 3$ by the recurrence
$P(n)=P(n-2)+P(n-3)\,.$ (6.8)
Hence we set $k=3$ and $(\beta_{1},\beta_{2},\beta_{3})=(0,1,1)$ in Sec. 6.3. Because $\beta_{1}=0$, we exclude all terms in $i_{1}$ in eq. (6.5) with $i_{1}>0$. This yields
$F(n+2)=\sum_{2i_{2}+3i_{3}=n}\binom{i_{2}+i_{3}}{i_{2},i_{3}}\,.$ (6.9)
With the initial values $F(0)=F(1)=0$, this yields the sequence $0,0,1,1,1,2,2,3,4,5,7,9,12,16,\dots$. Hence $P(n)=F(n+2)$.
6.6 Perrin sequence
The Perrin and Padovan sequences satisfy the same recurrence, but with different initial values. We write ‘$Q(n)$’ to avoid confusion with the Padovan sequence. Then $Q(0)=3$, $Q(1)=0$ and $Q(2)=2$ and for $n\geq 3$ the recurrence is
$Q(n)=Q(n-2)+Q(n-3)\,.$ (6.10)
Recall the sum in eq. (6.5) is normalized to yield $F(k-1)=F(2)=1$. Setting $F(0)=1$ and $F(1)=0$ yields
$Q(n)=2F(n)+P(n-5)=2P(n-2)+P(n-5)\,.$ (6.11)
Here $F(n)$ is given by eq. (6.9) and recall $P(n)$ is a Padovan number. For $n<5$, one must extend the Padovan sequence to negative arguments. The first few terms are $3,0,2,3,2,5,5,7,10,12,17,22,\dots$.
7 Generating function
The generating function for the $k$-generalized Fibonacci numbers is
$G(z)=\frac{z^{k-1}}{1-z-\dots-z^{k}}\,.$ (7.1)
Then $F_{n}(n)$ is the coefficient of $z^{n}$ in the series expansion of $G(z)$. Process the generating function as follows
$\begin{split}G(z)&=\frac{z^{k-1}}{1-z-\dots-z^{k}}\\ &=\frac{z^{k-1}}{1-z(1+\dots+z^{k-1})}\\ &=\frac{z^{k-1}}{1-z(1-z^{k})/(1-z)}\\ &=\frac{(1-z)z^{k-1}}{1-2z+z^{k+1}}\,.\end{split}$ (7.2)
Define $T=1/(1-2z+z^{k+1})$ and expand it in a series as follows
$\begin{split}T&=\frac{1}{1-2z+z^{k+1}}\\ &=\frac{1}{1-2z}\frac{1}{1+z^{k+1}/(1-2z)}\\ &=\frac{1}{1-2z}\sum_{s=0}^{\infty}(-1)^{s}\frac{z^{s(k+1)}}{(1-2z)^{s}}\\ &=\sum_{s=0}^{\infty}(-1)^{s}\frac{z^{s(k+1)}}{(1-2z)^{s+1}}\\ &=\sum_{s=0}^{\infty}(-1)^{s}z^{s(k+1)}\sum_{u=0}^{\infty}\binom{s+u}{s}(2z)^{u}\\ &=\sum_{j=0}^{\infty}z^{j}\sum_{s=0}^{\lfloor s/(k+1)\rfloor}(-1)^{s}2^{j-s(k+1)}\binom{j-sk}{s}\qquad(j=s(k+1)+u)\,.\end{split}$ (7.3)
Then $G(z)$ consists of a sum of two series
$\begin{split}G(z)&=z^{k-1}\sum_{j=0}^{\infty}z^{j}\sum_{s=0}^{\lfloor j/(k+1)\rfloor}(-1)^{s}2^{j-s(k+1)}\binom{j-sk}{s}\\ &\quad-z^{k}\sum_{j=0}^{\infty}z^{j}\sum_{s=0}^{\lfloor j/(k+1)\rfloor}(-1)^{s}2^{j-s(k+1)}\binom{j-sk}{s}\,.\end{split}$ (7.4)
The coefficient of $z^{n}$ is $F_{k}(n)$:
$\begin{split}F_{k}(n)&=\sum_{s=0}^{\lfloor(n-k+1)/(k+1)\rfloor}(-1)^{s}2^{n-k+1-s(k+1)}\binom{n-k(1+s)+1}{s}\\ &\quad-\sum_{s=0}^{\lfloor(n-k)/(k+1)\rfloor}(-1)^{s}2^{n-k-s(k+1)}\binom{n-k(1+s)}{s}\,.\end{split}$ (7.5)
We can set both upper limits equal to those in the first sum. (This may introduce an extraneous term in the second sum, but if so, it equals zero.) The final result is more elegant.
$F_{k}(n)=\sum_{s=0}^{\lfloor(n-k+1)/(k+1)\rfloor}(-1)^{s}2^{n-k-s(k+1)}\biggl{[}2\binom{n-k(1+s)+1}{s}-\binom{n-k(1+s)}{s}\biggr{]}\,.$ (7.6)
Observe that the sum is not nested, hence the computational complexity to evaluate $F_{k}(n)$ does not increase as the value of $k$ increases.
Remark 7.1.
A similar expression was obtained by Ferguson [12], where the series expansion of $T$ in eq. (7.3) was derived differently. See also Kessler and Schiff ([13] eq. (6)), which generalized Ferguson’s work beyond $k$-generalized Fibonacci numbers.
8 Basis sequences
Let us return to the Vandermonde system of equations in eq. (3.1). The $k$-generalized Fibonacci numbers are the solution for only one set of initial conditions, viz. $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. Let us derive a set of $k$ ‘basis sequences’ $B_{km}(n)$, $m=0,\dots,k-1$ as follows. First, for consistency of notation, we reexpress the recurrence eq. (1.1) as
$B_{k}(n)=B_{k}(n-1)+\dots+B_{k}(n-k)\,.$ (8.1)
The basis sequence $B_{km}(n)$ is the solution of eq. (8.1) for the the initial values $B_{km}(n)=\delta_{mn}$ for $10\leq m,n\leq k-$ 1. (The $k$-generalized Fibonacci numbers are the last member of the basis, viz. $m=k-1$, so $F_{k}(n)=B_{k,k-1}(n)$.) Then, for arbitrary initial values $(\gamma_{0},\dots,\gamma_{k-1})$ for $n=0,\dots,k-1$, the solution of the recurrence eq. (8.1) is
$B_{k}(n)=\sum_{m=0}^{k-1}\gamma_{m}B_{km}(n)\,.$ (8.2)
To proceed, we extend the Vandermonde system of equations as follows (attaching row and column indices to the coefficients $c$)
$\begin{split}\begin{pmatrix}1&1&\dots&1\\ \zeta_{0}&\zeta_{1}&\dots&\zeta_{k-1}\\ \zeta_{0}^{2}&\zeta_{1}^{2}&\dots&\zeta_{k-1}^{2}\\ \vdots&&&\vdots\\ \zeta_{0}^{k-1}&\zeta_{1}^{k-1}&\dots&\zeta_{k-1}^{k-1}\end{pmatrix}\begin{pmatrix}c_{00}&c_{01}&\dots&c_{0,k-1}\\ c_{10}&c_{11}&\dots&c_{1,k-1}\\ c_{20}&c_{21}&\dots&c_{2,k-1}\\ \vdots&&&\vdots\\ c_{k-1,0}&c_{k-1,1}&\dots&c_{k-1,k-1}\end{pmatrix}=\begin{pmatrix}1&0&\dots&0\\ 0&1&\dots&0\\ \vdots&\ddots&&\vdots\\ \vdots&&\ddots&\vdots\\ 0&0&\dots&1\end{pmatrix}\,.\end{split}$ (8.3)
The basis sequence $B_{km}(n)$ is given by
$B_{km}(n)=\sum_{\ell=0}^{k-1}c_{\ell m}\zeta_{\ell}^{n}\,.$ (8.4)
To determine $c_{\ell m}$, we invert the Vandermonde matrix in eq. (8.3). The inverse is the matrix whose rows are the coefficients of the Lagrange polynomials $L_{\ell}(x)$, $\ell=0,\dots,k-1$. The solution is $c_{\ell m}=L_{\ell m}$, which is the coefficient of $x^{m}$ in $L_{\ell}(x)$. Recall from eq. (5.1) the Lagrange polynomial $L_{\ell}(x)$ is
$L_{\ell}(x)=\frac{(x-\zeta_{0})\dots(x-\zeta_{\ell-1})(x-\zeta_{\ell+1})\dots(x-\zeta_{k-1})}{(\zeta_{\ell}-\zeta_{0})\dots(\zeta_{\ell}-\zeta_{\ell-1})(\zeta_{\ell}-\zeta_{\ell+1})\dots(\zeta_{\ell}-\zeta_{k-1})}\,.$ (8.5)
We denote the numerator by $N_{\ell}(x)$ and write $N_{\ell}(x)=\sum_{m=0}^{k-1}N_{\ell m}x^{m}$. The coefficients $N_{\ell m}$ have the following structure
  1. 1.
    First, $N_{\ell,k-1}$ equals unity.
  2. 2.
    Next, for $j=1,\dots,k-1$, $N_{\ell,k-1-j}=(-1)^{j}\times$ (sum of the products of $j$-tuples of roots excluding $\zeta_{\ell}$).
  3. 3.
    Hence $N_{\ell,k-2}$ equals the negative of the (sum of the roots excluding $\zeta_{\ell}$) and $N_{\ell,k-3}$ equals the sum of the (products of pairs of roots excluding $\zeta_{\ell}$), etc.
The coefficients $N_{\ell m}$ can be obtained from the auxiliary polynomial $A(x)$ as follows. First, define the sums of products of $j$-tuples of roots as follows $S_{j}=\sum_{0\leq i_{1}<\dots<i_{j}\leq k-1}\zeta_{i_{1}}\dots\zeta_{i_{j}}$, $j=1,\dots,k-1$. Recall $A(x)=x^{k}-x^{k-1}-\dots-1$ and by definition $A(x)=(x-\zeta_{0})\dots(x-\zeta_{k-1})$. From the theory of polynomials, observe that $S_{j}=(-1)^{j-1}$. Then some careful bookkeeping yields
$N_{\ell,k-1}$ $=1\,,$ (8.6a)
$N_{\ell,k-2}$ $=\zeta_{\ell}N_{\ell,k-1}-1$ $=$ $\;\zeta_{\ell}-1\,,$ (8.6b)
$N_{\ell,k-3}$ $=\zeta_{\ell}N_{\ell,k-2}-1$ $=$ $\;\zeta_{\ell}^{2}-\zeta_{\ell}-1\,.$ (8.6c)
In general, $N_{\ell,k-1}=1$ and $N_{\ell,m}=\zeta_{\ell}N_{\ell,m+1}-1$ and we work recursively to deduce, for $m=k-2,k-3,\dots,0$,
$N_{\ell,m}=\zeta_{\ell}^{k-m-1}-\zeta_{\ell}^{k-m-2}-\dots-1\,.$ (8.7)
The expression for the denominator $D_{\ell}$ of the Lagrange polynomial $L_{\ell}(x)$ was derived in eq. (5.4). Then the coefficient $c_{\ell m}$ is given by
$c_{\ell m}=N_{\ell m}\,\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\,\zeta_{\ell}^{1-k}\,.$ (8.8)
Finally, the basis sequence $B_{km}(n)$ is given by ($m=0,\dots,k-1$)
$B_{km}(n)=\sum_{\ell=0}^{k-1}N_{\ell m}\,\frac{\zeta_{\ell}-1}{(k+1)\zeta_{\ell}-2k}\,\zeta_{\ell}^{n-k+1}\,.$ (8.9)
This furnishes a basis (or ‘fundamental solutions’) for the recurrence eq. (8.1).
Note the following. The first and last basis sequences $B_{k0}(n)$ and $B_{k,k-1}(n)$ are related via a shift of the index $n$ as follows
$B_{k0}(n)=B_{k,k-1}(n-1)\,.$ (8.10)
Proof.
First note that $B_{k0}(k)=1$. Then observe that $\{B_{k0}(1),\dots,B_{k0}(k-1),B_{k0}(k)\}=\{0,\dots,0,1\}$. These are exactly the initial values for the sequence $B_{k,k-1}(n)$, but the index $n$ is shifted by unity. Eq. (8.10) follows immediately. ∎
As an example, let us apply the above results to the Fibonacci numbers, i.e. $k=2$. Denoting the Fibonacci numbers by $F_{n}$ and using eq. (8.10) yields $B_{2,0}(n)=F_{n-1}$ and $B_{2,1}(n)=F_{n}$. Then for the initial values $(\gamma_{0},\gamma_{1})$, the general solution is $B_{2}(n)=\gamma_{0}F_{n-1}+\gamma_{1}F_{n}$. This proves (yet again) that every solution of the standard Fibonacci recurrence can be expressed as a linear combination of $F_{n-1}$ and $F_{n}$. The Lucas numbers $L_{n}$ have the initial values $(L_{0},L_{1})=(2,1)$, whence
$L_{n}=2F_{n-1}+F_{n}=2F_{n+1}-F_{n}\,.$ (8.11)
This is the known relation of the Lucas numbers to the Fibonacci numbers.